Thanks for your continued help.
Can you Solve the following 2 equation:
x^2e^x-5xe^x=0
5^3x=5^4x-2

Answer by stanbon(57222)

Can you Solve the following 2 equation:
x^2e^x-5xe^x=0
Factor to get:
xe^x(x-5)=0
x=0 or e^x=0 or x-5=0
x=0 or x=5
Note e^x is never zero
----------------
5^3x=5^(4x-2)
3x=4x-2
x=2
-------------
Cheers,
Stan H.